Norman Margolus Tommaso Toffoli

STEP: a SpaceTime Event Processor

Data Sheets

(version 2.2.2)

MIT Laboratory for Computer Science Cambridge, MA 02139

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993, 1994, 1995 by Norman Margolus and Tommaso Toffoli

ii

STEP Data Sheets

/fab/ref/data/_pref

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 T. Toffoli DRAFT, February 2, 1995

Preface

This manual gathers in one place data sheets for offthe-shelf commercial components (semiconductors, connectors and sockets, mechanical hardware, etc.) used in a step unit. Custom components are described in the Fabrication Reference.

The data sheets are provided as a courtesy to step users, developers, and documenters, as well as to the components’ manufacturers. Where the manufaturer’s consent was not explicitly obtained, we presume implicit consent and encouragement for this policy. Most of the sheets are culled from data books bearing copyrights and disclaimers; distribution and use of the data sheets is subject to the understanding that such copyrights and disclaimers remain in force even if when they do not appear on the individual data sheets.

The data sheets appear, as listed in the Contents, in alphabetical order by manufacturer and by part number under each manufacturer. For convenience, manufacturer and part number are repeated as, respectively, major and minor header on each page; page numbering starts afresh with each part.

Thus, the header

Motorola MC88915 [3]

denotes page 3 of the data sheets for the device MC88915 made by Motorola.

STEP Data Sheets

/fab/ref/data/man.toc

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 T. Toffoli DRAFT, February 2, 1995

Contents

AMP

Contact Arrangements
Panel Cutout Specs

205817-3 (VGA screwlock set) 748565-1 (VGA panel receptacle) 748610-4 (VGA female pins) 749075-5 (right-angle header) 749076-5 (right-angle header) 749084-5 (right-angle plug)

749086-1/087-1/087-2 (screwlock kits)

749611-5 (panel receptacle) 749621-5 (cable plug) 749649-5 (right-angle header) 749656-5 (panel receptacle) 749699-5 (cable receptacle) 749877-5 (panel receptacle) 821977-1 (32-pin PLCC socket)

Atmel

AT27C256R (256K EPROM) Augat

PCS-020/028SMU-11 (20/28-pin PLCC sockets) Berg

65043-032 (mini-latch housing) 47748-000 (crimp receptacle) 66245-250 (backplane receptacle) 66247-008 (backplane shroud)

Brooktree Bt478 (color palette RAMDAC) Bt858 (RGB-to-video encoder)

Data Delay Devices

DDU-7F-10 (digital delay unit) Fujitsu

FCN-234P096-G/0 (SBus connector)

Hirose

HR10A-10R-12S (circular panel mount recept.) IDT

IDT72211 (FIFO)

Lattice Semiconductor GAL16V8B (programmable logic device) GAL22V10B (programmable logic device)

LZRelectronics

LZR-MDJCP4 (mini-DIN panel mount recept.) Maxim

MAX698 (power-on reset) Molex

03-06-1023/2023 (recept. /plug shell) 02-06-1103/2103 (female/male pin)

Motorola MC10H645 (clock buffer) MC88915 (phase-locked loop)

MUPAC

3351050-02-PKG (inserter/ejector) National Semiconductor BAV99 (silicon diode pair) LM385-1.2 (voltage reference) 74AC08 (quad AND gate) 74AC377 (octal D flip-flop) 74ACQ244 (octal buffer) 74ACQ245 (octal transceiver) 74ACQ374 (octal D flip-flop)

OKI Semiconductor

                                                 MSM514402 (4M DRAM)

Texas Instruments

TIBPAL16L8/16R4 (programmable array logic) Toshiba

                                                 TC5117400 (16M DRAM)
                                                 TC514400 (4M DRAM)

TC524256 (1M video RAM) TC55464 (256K SRAM)

TRW TMC1175 (A/D converter)

[1] 74AC08 National Semiconductor National Semiconductor 74AC08 [2] [3] 74AC08 National Semiconductor National Semiconductor 74AC08 [4]